E907 Systems Conference Call Minutes

7 November 2002

Location

Phone Number

Access Code

WH-11SW

805-240-9463

536495

Contents

Progress Updates and Issues

  1. Upstream Beamline
    1. BEAM - Beamline
    2. UBL - Upstream Beamline Detectors
  2. Targets
    1. TGTW - Target Wheel
    2. CTGT - Cryo Targets
    3. NTGT - NuMI (MINOS) Target
  3. TPC - Time Projection Chamber
  4. Magnets
    1. ZipTrack
    2. JGG - Jolly Green Giant
    3. Rosy
  5. CKOV - Cerenkov
  6. DC - Chambers
    1. DC1-4 (E690 Chambers)
    2. DC5 & DC6 (Iowa Chambers)
  7. TOF - Time of Flight
  8. RICH - Ring Imaging Cerenkov
  9. ECAL - EM Calorimeter
  10. NCAL - Hadron Calorimeter
  11. Gas Systems
  12. DAQ - Data Acquisition
  13. MC - Monte Carlo
  14. Installation
    1. Alignment
  15. MC7 - Enclosure and Counting House
  16. Meetings
  17. Project Management

Next Call


Progress Updates and Issues

Leon Beverly's meeting summaries in italics.

Upstream Beamline

BEAM - Beamline

Raja:

We need to meet to discuss the design of the MIPP secondary beam.

Time:- 5:00PM, Friday Nov 8,2002

Place:- WH12E conference room.

Topics:-

  1. Dispersion at the momentum slit. Is it too small? How do we fix the design
  2. When we tune to 90 GeV positives, How many primary protons punch through? How do we ameliorate this problem?
  3. Is it possible to get a smaller angular divergence at the secondary target?
Dave Cary reports that he still looking at all aspects of the beam, final focus transport, etc. Based upon unofficial comment, FY03 funding for the Meson 120 program has the support of the Beams Division leader.

UBL - Upstream Beamline Detectors

BCKOV

Need to locate a reflectrometer to get a before and after resurfacing number for the mirrors. Jim will send out one mirror to have the center made larger, have it evaluated for stripping and resurfacing. After the experience, Livermore will place a PO for the balance of the mirror resurfacing.

Targets

TGTW - Target Wheel

CTGT - Cryogenic Targets

NTGT - NuMI (MINOS) Target

TPC - Time Projection Chamber

Bob Jones group has an operating clock module, powered up and operating stick and will proceed to trouble shoot non-operational sticks by comparing signals to the operating stick.

Magnets

Main LCW header installation, which is needed to ziptrack, - John C. will ok the T&M fitter requisition.

ZipTrack

Walt is working with Holger Meyer to get hall probes mounted in the ANs.

Some shielding of ziptrack hall probes is ongoing at lab 8.

JGG - Jolly Green Giant

The coil cooling hoses and Thermal switches (klixons) have been connected. Inner coil DC leads will be installed on JGG the week of 11/11.

Rosie

DS side of Rosie coil shims installation started the week of 11/4. That work will be followed by connect LCW hoses. Fabrication of Bus jumpers to series coils is in process. Klixons need to be connected.

CKOV - Cerenkov

Move from PC4 to MC7 about the last week in November (may be later if ziptracking is not completed).

DC - Drift Chambers

DC1-4 (E690 Chambers)

Extensive discussion with Nick Solomay. Bottom line, chambers “look good,” but further tests need to be completed.  

DC#3 support stand drawing complete & ready for checking. DC#4 stand drawing next.

DC5 & DC6 (Iowa Chambers)

A recycler from KTeV may be used to allow circulation of Freon or dimethylether through the Iowa chambers. Terry will also look into the recycler status. Nick will also look at the RMH electronics.

The support stand is ready for fabrication quote once drawings are signed off by Ed LaVallie.

Nick Solomey:

  1. Discussed with Karen Kephart for an hour friday afternoon and have some recomendations regarding the small wire chambers.
  2. Talked with John Krider about the TRD KTeV recirculator we will need to change two parts to make it work with Freon. Small price if we can use them.
  3. Had an e-mail exchange with Yasar at Iowa and have several comments about the larger wire chambers.

Pierrick Hanlet:

Saturday was a fairly successful day. We did not look at any signals in the chambers, but we did take apart a chamber that we have previously opened. In doing this, we learned how to safely perform this task without additional heavy reliance on Ed.

There has been some question about the tension of the cathode wires. We again measured them, as well as the anode wires and found that the cathode wires have between 90--110g tension. The anode wires are at about 50g tension. These values were consistent for every measurement that we made. Adam performed 3 calculations based on different references; unfortunately, these were not consistent. We chose to believe the formula in Sauli, which depends on the capacitance on the wires. Using reasonable values of capacitance, we found that both the anode tension and cathode tension were both theoretically reasonable for our chambers, though they are close to the lower limit. While the wires may have actually been wound at 150g for cathode and 90g for anode, it is not known how much the tension was reduced when the frames were transferred from the winding machine to the chamber.

We found quite a bit of dirt in the chambers. Other than for one wire, we did not find any strong correlation between Adam's previous measurements of high capacitance and individual wires. Unfortunately, since we did not know how the chambers were constructed, Adam's capacitance measurements were made between the wire and a floating ground. (Adam, please correct me if I've got this wrong.)

We also measured the ground planes resistance to the frame. For one plane, we found that the wires have the same 0.042" spacing as the cathode wires and each wire has a resistance of 20Mohms to ground. The other ground plane has a wire spacing that appears to be about doubled the spacing of the cathode planes, and has a resistance of 10Mohms to ground. Here we found a few channels were shorted to ground and a few others had an open circuit to ground. The shorts are not a problem unless there are sparks near those wires, but we should have connections to ground for the open wires. The wires with the open circuit appear to have a kludge which gave way.

The plan now is to look at signals with a source on the working chamber. Assuming that all wire planes have similar tensions as those we just measured, we will use real signals to verify that the tensions are adequate a la suggestions of Ken. We will do this this week and report on it this weekend.

At this point, the plan for the open chamber is to rebuild it after acquiring new gasket material and finishing the cleaning (which is not much). With each anode plane that is put down, I believe that we should check for continuity for each wire; any open circuit should be repaired. The open wires on the ground plane should be repaired. Once the chamber is buttoned up, we should improve the grounding of both sides of each connector card to the window frames, and add some grounding straps between the two window frames along the dimension normal to the electronics. We may also consider ways in which to improve the HV distribution.

If we find the first chamber to be ok, we will apply the grounding improvements and give the chamber to Dave Northacker to install the electronics mounting. We will then move the second chamber to the testing area, fill it with working gas and electronics, and test it with a source.

We will then rotate the large chamber onto the dismantling table and repeat what we did to the presently opened chamber, etc. etc.

I think that we are all pretty happy with the results from Saturday. We might get lucky, and won't have to rewind anything.

On my part, I will continue with the electronics testing. Thanks to Bill Luebke and Walter Jaskierny, we have made great progress in getting a working system. I had a dozen preamps tested for Saturday, but we didn't need them. We will use them this week to test the chamber. Adam has requested some tested discriminator cards; I won't be able to provide those until next week as we have to turn our working prototypes into real boards.

Adam Bujak:

My original measurement of capacitances has been performed correctly: chambers intact, ground pins on all connectors grounded, also - all wires grounded, except the wires just measured. We made a mistake this Saturday trying to find correlation with previous measurements on disassembled planes: this made no sense since there was no ground. Based on my experience from CMS I'm sure my results are correct.

Independently from our short term goals we should remember what is our ultimate goal, which is to make chambers work. I do not excpect conclusive results from a check on randomly chosen wires (note restricted access to the window). In my opinion the correct course of action is as follows:

  1. have a complete set of good (tested) preamps, discriminators and cables;
  2. fully instrument a chamber (grounding, external inteference and noise conditions change with a number of cables connected);
  3. take cosmics and determine an overall characteristics of a chamber, and only then
  4. investigate trouble spots with a source.

What kept me (since almost 3 months ago) from proceeding in this manner with the chamber no.1 were bad preamp/discriminator/tdc channels. Recently another two discriminator boards in my setup have died. There are low efficiency regions in the chamber, but it's difficult to judge what is to blame - electronics or the chamber itself. Random check with a source will not answer the question. I think, that speeding up the electronics testing should be our highest priority.

TOF - Time of Flight

RICH - Ring Imaging Cerenkov

14th floor has front-end board (32 channel) working via a debug port, but not through a VME crate yet.

ECAL - EM Calorimeter

Mike Longo:

We are happy to share the responsibility for the hadron calorimeter (HCAL) with the Univ. of Virginia group.

Since they built the HCAL for HyperCP, a reasonable division of responsibility is for us to be responsible for the mechanical modifications required for MIPP. In particular, we would contribute up to $5000 for the new base for the existing calorimeter stand. The Univ. of Virginia group would be responsible for the phototubes, light-pulser calibration system, and readout electronics.

Responsibility for testing, calibration, and data analysis would be shared by the two groups.

The Univ. of Michigan group would have complete responsibility for the EMCAL, including construction, testing, and electronics. We will move forward with its design and construction in conjunction with Fermilab engineers.

We expect the Academic Alliances money to "arrive" around Nov. 15. I'll set up the FNAL account as soon as I can after that. We may need to 'borrow' a bit of Ingrid's time before that happens.

After HA & lead training is complete, assembly may start about 12/1.

NCAL - Calorimeter

Gas Systems

DAQ - Data Acquisition

David Asner:

We will be having a DAQ meeting at 1p-3p this friday. Same room as last time. The conference room on the third floor of the FCC. Agenda to follow

MC - Monte Carlo

Installation

Alignment

The set up of a survey network using the laser tracker is nearly complete.

MC7 Enclosure and Counting House

Installation conduit and wiring to power heaters, OH doors, exit & EM lights in progress.

Meetings

Project Management

Next Systems Call

The next call will be Thursday, 14 November 2002, at 3:00 CDT in WH-3NE, Theory Room.


PDB, 12/10/02